### v2

@@ Abstract
Scene Graph Generation (SGG) aims to detect all the visual relation triplets $<$sub, pred, obj$>$ in a given image. With the emergence of various advanced techniques for better utilizing both the intrinsic and extrinsic information in each relation triplet, SGG has achieved great progress over the recent years. However, due to the ubiquitous long-tailed predicate distributions, today’s SGG models are still easily biased to the head predicates. Currently, the most prevalent debiasing solutions for SGG are re-balancing methods, e.g., changing the distributions of original training samples. In this paper, we argue that all existing re-balancing strategies fail to increase the diversity of the relation triplet features of each predicate, which is critical for robust SGG. To this end, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy, which aims to mitigate the bias issue from the perspective of increasing the diversity of triplet features. Specifically, we enhance relationship diversity on not only feature-level, i.e., replacing the intrinsic or extrinsic visual features of triplets with other correlated samples to create novel feature compositions for tail predicates, but also image-level, i.e., manipulating the image to generate brand new visual appearance for triplets.
Due to its model-agnostic nature, MCFA can be seamlessly incorporated into various SGG frameworks. Extensive ablations have shown that MCFA achieves a new state-of-the-art performance on the trade-off between different metrics.
Keywords:
Unbiased Scene Graph Generation Feature Augmentation Category ClusteringCompositional Learning
@@ 1 Introduction
As one of the fundamental comprehensive visual scene understanding tasks, Scene Graph Generation (SGG) has attracted unprecedented interest from our community and has made great progress in recent years Lu et al. (2016); Zhang et al. (2017); Zellers et al. (2018); Chen et al. (2019a, b); Tang et al. (2020); Li et al. (2021b, 2022a). Specifically, SGG aims to transform an image into a visually-grounded graph representation (i.e., scene graph) where each node represents an object instance with a bounding box and each directed edge represents the corresponding predicate between the two objects. Thus, each scene graph can also be formulated as a set of visual relation triplets (i.e., $<$sub, pred, obj$>$). Since such structural representations can provide strong explainable potentials, SGG has been widely-used in various downstream tasks, such as visual question answering, image retrieval, and captioning.
@@
Figure 1: (a) The intrinsic and extrinsic information for SGG. The entity prediction is for the green box, and the predicate prediction is for the relation between the red and green boxes. (b) Illustration of the diversity of feature space and decision boundary between on and laying on before and after using re-balancing and MCFA. Each sample denotes the corresponding visual triplet features.
In general, due to the extremely diverse visual appearance of different visual relation triplets, recent SGG methods all consider both intrinsic and extrinsic information for entity and predicate classification Zellers et al. (2018); Tang et al. (2019); Lu et al. (2021).
@@
By “intrinsic information”, we mean these intrinsic characteristics of the subjects and objects, such as their visual, semantic, and spatial features. For example in Fig. 1(a), with the help of these intrinsic features, we can easily infer all possible entity categories (e.g., pole or surfboard) and predicate categories (e.g., on or holding) of each triplet. However, sometimes it is still hard to confirm the exact correct predictions with only intrinsic information, especially for tiny objects. Thus, it is also essential to consider other “extrinsic information” in the same image, such as the context features from neighbor objects. As shown in Fig. 1(a), after encoding the features of surrounding objects (e.g., wave and hand), we can easily infer that the categories of the entity and predicate should be surfboard and holding.
@@
Although numerous advanced techniques have been proposed to enhance SGG performance. Today’s SGG methods still fail to predict some informative predicates due to the ubiquitous long-tailed predicate distribution in prevalent SGG datasets Krishna et al. (2017). Such a distribution is characterized by a small amount categories with vast samples (head [^1] We directly use “tail”, “body”, and “head” categories to represent the predicate categories in the tail, body, and head parts of the number distributions of different predicates in SGG datasets, respectively.) and plenty of categories with rare samples (tail). As a result of the discrepancy of sample size among different categories, the learned decision boundary becomes improper (cf.. Fig. 1(b)), i.e., their predictions are biased towards the head predicates (e.g., on) and they are error-prone for the tail ones (e.g., laying on).
@@
To overcome the bias issue, the most prevalent unbiased SGG solutions are re-balancing strategies, e.g., sample re-sampling Li et al. (2021b) and loss re-weighting Yan et al. (2020); Chen et al. (2022); Li et al. (2022c); Lyu et al. (2022). They alleviate the negative impact of long-tailed distribution by increasing samples or loss weights of tail classes. Then, the decision boundaries are adjusted to reduce the bias introduced by imbalanced distributions. However, we argue that all the existing re-balancing strategies only change the frequencies or contributions of existing relation triplets to improve the skewed decision boundaries, but fail to enhance the diversity of relation triplet features [^2] We use the “relation triplet feature” to represent the combination of both intrinsic and extrinsic features of each visual relation triplet. of each predicate, which is crucial to infer their complete feature space. This sparsity of feature diversity will lead to under-representation of tail categories, thereby making it hard to infer the complete data distribution, i.e., making it challenging to find the optimal direction to adjust the decision boundaries Chu et al. (2020); Vigneswaran et al. (2021). For example in Fig. 1(b), the feature space of laying on is so sparse that the decision boundary can be adjusted within a large range. The performance of this naive adjustment without “complete” distribution is always sensitive to hyperparameters, i.e., excessively increasing the sample number or loss weight of tail predicates may cause some head predicate samples to be incorrectly predicted as tail classes (right solid line), and vice versa (left solid line).
@@
Figure 2: (a) Intrinsic-CFA: Replacing the entity feature of tail predicate triplet dog-laying on-bed from dog to cat to enhance the intrinsic feature. (b) Extrinsic-CFA: Mixing up the feature of tail predicate triplet pillow-laying on-bed into the context of pillow-on-bed to enhance the extrinsic feature.
Therefore, in this paper, we focus on enhancing diversity of relation triplet features instead of re-balancing their frequencies to learn more robust feature space and decision boundaries for predicates. One intuitive idea is to borrow novel entity features or context features from other existing triplet samples to replace the original features, thereby constructing new triplet features for the target predicates. By deeply exploring the correlations of objects and predicates, we can reasonably select the features to replace, ensuring the reliability of the re-composed relation triplet features. However, this naive augmentation technique on feature-level manipulation has its inevitable limitations. On the one hand, directly replacing the original visual features may lead to unnatural transitions (e.g., fusing a pillow’s feature with another one that are inadvertently entangled with adjacent objects, such as bedside cupboard) as shown in Fig. 3(b), further hurting the expressive ability of the model. On the other hand, the augmented samples are still limited to the in-domain distribution, and as a result, the trained models are unable to handle with out-of-domain predictions. Therefore, we also propose to manipulate relation triplet features on image-level with the help of powerful generative models as shown in Fig. 3(c).
Figure 3: (a) Context Image: The image which provides context information and is used to to augment with the triplet pillow-on-bed. (b) Feature-level Augmentation: Visualization of augmented results with the feature of tail predicate triplet pillow-laying on-bed selected from dataset. (c) Image-level Augmentation: Augmented new image with target triplet and context information by generative models.
@@
In this paper, we propose a novel Multi-level Compositional Feature Augmentation (MCFA) strategy for unbiased SGG, which tries to solve the bias issue by enhancing the diversity of relation triplet features on both feature-level and image-level, especially for tail predicates.
As shown in Fig. 1(b), by enhancing the feature diversity of the tail predicates (e.g., laying on), SGG models can easily learn the proper decision boundaries (vs. re-balancing strategies).
@@
On feature-level, we apply different strategies for augmenting intrinsic and extrinsic feature diversity respectively. For intrinsic diversity, we replace the entity features (e.g., subject or object) of a tail predicate triplet with other “suitable” entity features. To determine the suitable entity categories for the augmentation, we propose a novel relation-conditioned entity clustering method to find the predicate-aware correlations between different entity categories, and then we regard the entity features from the same cluster are suitable. Specifically, we calculate the category correlation by pattern, context, and semantic similarities. For example in Fig. 2 (a), the entity categories cat and dog are in the same cluster, i.e., we can augment the intrinsic feature by replacing the dog entity feature with a cat entity feature. For extrinsic diversity, we take advantage of the context or interactions of other triplets (i.e., context triplets) and enhance the features of tail predicate triplets by these context triplets. Specifically, given a context triplet randomly selected from an image, we first select a reasonable tail predicate triplet as the target by limiting the categories and relative position of two objects. Then, to minimize the impact on the prediction of other triplets in the original image and make use of the extrinsic features of the context triplet, we use an adaptive mixup strategy to fuse the features of targeted tail predicate triplet into the context triplet and adaptively adjust mixup weights based on predicate semantic similarity. For example in Fig. 2 (b), triplet pillow-laying on-bed is mixed up into the image of pillow-on-bed. To avoid introducing noise into augmented samples, we adaptively assign weights to the mixed triplets by considering the semantic similarity of relation triplets.
On image-level, we aim to directly manipulate the images and generate brand-new visual appearances for target predicate triplets. A core challenge in this process is ensuring the semantic plausibility and spatial coherence of the generated images. To address this, we first employ a relative position constraint to filter only contextually appropriate images associated with target triplet categories for augmentation. For the synthesis process, we leverage the powerful grounded text-to-image generation capabilities of recent generative models (e.g., GLIGEN Li et al. (2023b)), as shown in Fig. 3(c). Specifically, the model is conditioned on the original image, the target triplet’s categories, and corresponding entity grounding information to generate a new high-quality visual instance. Subsequently, we assess the quality of each generated image via a visual-language similarity score to ensure high fidelity and relevance. As the context triplets are selected from other predicate triplets and visual features of target triplets are novelly generated, both extrinsic and intrinsic diversity are enhanced under our image-level augmentation strategy.
@@
We evaluate MCFA on two most prevalent and challenging SGG datasets: Visual Genome (VGKrishna et al. (2017) and GQA Hudson and Manning (2019). Since MCFA is a model-agnostic debiasing strategy, it can be seamlessly incorporated into various SGG architectures and consistently improve their performance. Unsurprisingly, MCFA can achieve a new state-of-the-art performance on the trade-off between different metrics. Extensive ablations and results on multiple SGG tasks and backbones have shown the generalization ability and effectiveness of MCFA.
This paper is a comprehensive and in-depth extension of our previous ICCV work Li et al. (2023a) with several significant improvements:
@@
    We propose a novel relation-conditioned entity clustering strategy for feature-level intrinsic augmentation by leveraging the predicate-driven semantic knowledge, leading to more accurate groupings.
@@
    We propose an adaptive mix-up strategy for feature-level extrinsic augmentation, which dynamically adjust mix-up coefficient based on predicate similarity, effectively handling with the varying semantic context of triplets.
@@
    We devise a novel generative augmentation method to manipulate relation triplet features on the image-level and generate high-quality samples, thereby effectively boosting the performance of tail predicates.
  4. 4.
    Extensive qualitative and quantitative analyses demonstrate the effectiveness of our MCFA.
@@ 2 Related Work
Unbiased Scene Graph Generation. Biased predictions prevent further use of scene graphs in real-world applications. Recent unbiased SGG works can be divided into three main categories: 1) Re-balancing: It alleviates the negative impact of long-tailed predicate distribution by re-weighting or re-sampling Yan et al. (2020); Li et al. (2022c); Lyu et al. (2022); Li et al. (2021b); Desai et al. (2021). 2) Unbiased Inference: It makes unbiased predictions based on biased models Tang et al. (2020); Yu et al. (2021). 3) Noisy Label Learning: It reformulates SGG as a noisy label learning problem and corrects the noisy samples Li et al. (2022a, b). This work points out the drawbacks of existing re-balancing methods and studies unbiased SGG from the new perspective of multi-level feature augmentation.
Feature Augmentation. Data augmentation is a prevalent training trick to improve models’ performance. Conventional data augmentation methods DeVries and Taylor (2017); Zhong et al. (2020) usually synthesize new samples in a hand-crafted manner. Compared to these image-level data augmentation methods, feature augmentation is another efficient way to improve models’ generalizability by directly synthesizing samples in the feature space Chu et al. (2020); Li et al. (2021a). Compared to existing data augmentation applications, SGG is a sophisticated task that involves intrinsic and extrinsic features. In this work, we propose MCFA to enrich the diversity of relation triplet features for debiasing from both feature-level and image-level.
Compositional Learning (CL). CL has been successfully applied to various computer vision tasks. As for visual scene understanding, some Human Object Interaction (HOI) detection works Hou et al. (2020); Kato et al. (2018); Hou et al. (2021) compose new interaction samples that significantly benefit both low-shot and zero-shot settings. Inspired by these successes, we adopt this compositional approach for unbiased SGG. Our method enhances triplet feature diversity by first decomposing representations into their intrinsic and extrinsic constituents, and then augmenting these two feature streams.
Figure 4: The illustration of unbiased SGG framework with MCFA on feature-level augmentation.
Figure 5: The pipeline of feature-level and image-level augmentation strategy. Feature-level strategy consists of two approaches to generate new compositions of triplet features tailored for intrinsic and extrinsic diversity enhancement respectively, i.e., (a) Intrinsic-CFA and (b) Extrinsic-CFA. While image-level strategy generates new visual appearance of relation triplet by generative models to ensure the augmentation quality, i.e.(c) Image-CFA.
3 Multi-level Compositional Feature Augmentation
@@
In this section, we first revisit the two-stage SGG baselines in Sec. 3.1. Then, we elaborate the pipeline of MCFA, including feature-level and image-level feature augmentation strategy (cf., Fig. 4) in Sec. 3.2 and Sec. 3.3, respectively. Finally, we detail the training objectives in Sec. 3.4.
@@ 3.1 Revisiting the Two-Stage SGG Baselines
Since the mainstream SGG methods are two-stage models, we review the two-stage SGG framework here Zellers et al. (2018); Tang et al. (2019). A typical two-stage SGG model involves three steps: proposal generation, entity classification, and predicate classification. Thus, the SGG task $P(\mathcal{G}|I)$ is decomposed into:
@@
Proposal Generation $P(\mathcal{B}|I)$. This step aims to generate all bbox proposals $\mathcal{B}$. Given an image $I$, they first utilize an off-the-shelf object detector (e.g., Faster R-CNN Ren et al. (2015)) to detect all the proposals $\mathcal{B}$ and their visual features $\{\bm{v}_{i}\}$.
Entity Classification $P(\mathcal{O}|\mathcal{B},I)$. This step mainly predicts the entity category of each $\bm{b}_{i}$in $\mathcal{B}$. Given a visual feature $\bm{v}_{i}$ and proposal $\bm{b}_{i}$, they use an object context encoder $\mathtt{Enc}_{obj}$ to extract the contextual entity representation $\bm{f}_{i}$:
@@
where $\bm{w}_{i}$ is the GloVe embedding Pennington et al. (2014) of predicted $\widehat{o}_{i}$. It is worth noting that both $\mathtt{Enc}_{cls}$ and $\mathtt{Enc}_{rel}$ often adopt a sequence model (e.g., Bi-LSTM Zellers et al. (2018), Tree-LSTM Tang et al. (2019), or Transformer Yu et al. (2021)) to better capture context. After relation feature encoding, they use a relation classifier $\mathtt{Cls}_{rel}$ to predict the relation $\widehat{r}_{ij}$ between any two subject-object pairs:
@@
3.2 Feature-level CFA
In this paper, we treat a relation triplet feature2 in SGG as a combination of two components: intrinsic feature and extrinsic feature, which refer to the intrinsic information (i.e., the subject and object itself) and extrinsic information (i.e., the contextual objects and stuff), respectively. Correspondingly, our feature-level CFA consists of intrinsic-CFA and extrinsic-CFA to augment these two types of features. For ease of presentation, we call the “targeted tail predicate triplet whose feature to be enhanced” as “query triplet”, and the “relation triplet whose image is used to provide context” as “context triplet”. Besides, to facilitate the feature augmentation, we store all the visual features of all tail predicate triplets (i.e., visual features $\bm{v}$ of two entities and their union feature $\bm{u}$) in a feature bank before training. Besides, we adopt the repeat factor $\eta$ = $\max(1,\eta_{r})$ Gupta et al. (2019); Li et al. (2021b) to sample images to provide enough context triplet and query triplet for augmentation, where $\eta_{r}$ = $\sqrt{\lambda/{f_{r}}}$, $f_{r}$ is the frequency of predicate category $r$ on the entire dataset, and $\lambda$ is hyperparameter. The unbiased SGG pipeline with feature-level CFA is shown in Fig. 4.

### v3

@@ Abstract
Scene Graph Generation (SGG) aims to detect all the visual relation triplets <<sub, pred, obj>> in a given image. With the emergence of various advanced techniques for better utilizing both the intrinsic and extrinsic information in each relation triplet, SGG has achieved great progress over the recent years. However, due to the ubiquitous long-tailed predicate distributions, today’s SGG models are still easily biased to the head predicates. Currently, the most prevalent debiasing solutions for SGG are re-balancing methods, e.g., changing the distributions of original training samples. In this paper, we argue that all existing re-balancing strategies fail to increase the diversity of the relation triplet features of each predicate, which is critical for robust SGG. To this end, we propose a novel Compositional Feature Augmentation (CFA) strategy, which is the first unbiased SGG work to mitigate the bias issue from the perspective of increasing the diversity of triplet features. Specifically, we first decompose each relation triplet feature into two components: intrinsic feature and extrinsic feature, which correspond to the intrinsic characteristics and extrinsic contexts of a relation triplet, respectively. Then, we design two different feature augmentation modules to enrich the feature diversity of original relation triplets by replacing or mixing up either their intrinsic or extrinsic features from other samples. Due to its model-agnostic nature, CFA can be seamlessly incorporated into various SGG frameworks. Extensive ablations have shown that CFA achieves a new state-of-the-art performance on the trade-off between different metrics.
Corresponding author. Work was done when Lin Li visited HKUST.
@@ 1 Introduction
As one of the fundamental comprehensive visual scene understanding tasks, Scene Graph Generation (SGG) has attracted unprecedented interest from our community and has made great progress in recent years [29, 46, 44, 2, 3, 35, 24, 26, 21, 43, 23, 32]. Specifically, SGG aims to transform an image into a visually-grounded graph representation (i.e., scene graph) where each node represents an object instance with a bounding box and each directed edge represents the corresponding predicate between the two objects. Thus, each scene graph can also be formulated as a set of visual relation triplets (i.e., <<sub, pred, obj>>). Since such structural representations can provide strong explainable potentials, SGG has been widely-used in various downstream tasks, such as visual question answering, image retrieval, and captioning.
@@
Figure 1: (a) The intrinsic and extrinsic information for SGG. The entity prediction is for the green box, and the predicate prediction is for the relation between the red and green boxes. (b) Illustration of the diversity of feature space and decision boundary between on and laying on before and after using re-balancing and CFA. Each sample denotes the corresponding visual triplet features.
@@
In general, due to the extremely diverse visual appearance of different visual relation triplets, recent SGG methods all consider both intrinsic and extrinsic information for entity and predicate classification [44, 36, 30]. By “intrinsic information”, we mean these intrinsic characteristics of the subjects and objects, such as their visual, semantic, and spatial features. For example in Figure 1(a), with the help of these intrinsic features, we can easily infer all possible entity categories (e.g., pole or surfboard) and predicate categories (e.g., on or holding) of each triplet. However, sometimes it is still hard to confirm the exact correct predictions with only intrinsic information, especially for tiny objects. Thus, it is also essential to consider other “extrinsic information” in the same image, such as the context features from neighbor objects. As shown in Figure 1(a), after encoding the features of surrounding objects (e.g., wave and hand), we can easily infer that the categories of the entity and predicate should be surfboard and holding.
@@
Although numerous advanced techniques have been proposed to effectively leverage both intrinsic and extrinsic information, today’s SGG methods still fail to predict some informative predicates due to the ubiquitous long-tailed predicate distribution in prevalent SGG datasets [20]. Such a distribution is characterized by few categories with vast samples (head [^1] We directly use “tail”, “body”, and “head” categories to represent the predicate categories in the tail, body, and head parts of the number distributions of different predicates in SGG datasets, respectively.) and many categories with rare samples (tail). Since the discrepancy of feature diversity and sample size among different categories, the learned decision boundary becomes improper (c.f. Figure 1(b)), i.e., their predictions are biased towards the head predicates (e.g., on) and they are error-prone for the tail ones (e.g., laying on).
@@
To overcome the bias issue, the most prevalent unbiased SGG solutions are re-balancing strategies, e.g., sample re-sampling [26, 49] and loss re-weighting [41, 1, 27, 31, 17]. They alleviate the negative impact of long-tailed distribution by increasing samples or loss weights of tail classes. Then, the decision boundaries are adjusted to reduce the bias introduced by imbalanced distributions. However, we argue that all the existing re-balancing strategies fail to increase the diversity of relation triplet features [^2] We use the “relation triplet feature” to represent the combination of both intrinsic and extrinsic features of each visual relation triplet. of each predicate, i.e., they only change the frequencies or contributions of existing relation triplet features (c.f. Figure 1(b)). Since these tail categories are under-represented, it is still hard to infer the complete data distribution, i.e., making it challenging to find the optimal direction to adjust the decision boundaries [6, 38]. For example in Figure 1(b), the feature space of laying on is so sparse that the decision boundary can be adjusted within a large range. The performance of this naive adjustment without “complete” distribution is always sensitive to hyperparameters, i.e., excessively increasing the sample number or loss weight of tail predicates may cause some head predicate samples to be incorrectly predicted as tail classes (right solid line), and vice versa (left solid line).
@@
Figure 2: (a) Intrinsic-CFA: Replacing the entity feature of tail predicate triplet dog-laying on-bed from dog to cat to enhance the intrinsic feature. (b) Extrnisc-CFA: Mixing up the feature of tail predicate triplet pillow-laying on-bed into the context of pillow-on-bed to enhance the extrinsic feature.
@@
In this paper, we propose a novel Compositional Feature Augmentation (CFA) strategy for unbiased SGG, which tries to solve the bias issue by enhancing the diversity of relation triplet features, especially for tail predicates. Specifically, CFA consists of two components: an intrinsic feature augmentation (intrinsic-CFA) and an extrinsic feature augmentation (extrinsic-CFA), which enhance the intrinsic and extrinsic features2, respectively. As shown in Figure 1(b), by increasing the feature diversity of the tail predicates (e.g., laying on), SGG models can easily learn the proper decision boundaries (vs. re-balancing strategies).
@@
For intrinsic-CFA, we replace the entity features (e.g., subject or object) of a tail predicate triplet with other “suitable” entity features. To determine the suitable entity categories for the augmentation, we propose a new hierarchical clustering method to find the correlations between different entity categories, and then we regard the entity features from the same cluster are suitable. Specifically, we calculate the category correlation by pattern, context, and semantic similarities. For example in Figure 2(a), the entity categories cat and dog are in the same cluster, i.e., we can augment the intrinsic feature by replacing the dog entity feature with a cat entity feature. For extrinsic-CFA, we take advantage of the context or interactions of other triplets (i.e., context triplets) and enhance the features of tail predicate triplets by these context triplets. Specifically, given a context triplet randomly selected from an image, we first select a reasonable tail predicate triplet as the target by limiting the categories and relative position of two objects. Then, to minimize the impact on the prediction of other triplets in the original image and make use of the extrinsic features of the context triplet, we use a mixup operation to fuse the features of targeted tail predicate triplet into the context triplet. For example in Figure 2(b), triplet pillow-laying on-bed is mixed up into the image of pillow-on-bed.
@@
We evaluate CFA on two most prevalent and challenging SGG datasets: Visual Genome (VG[20] and GQA [16]. Since CFA is a model-agnostic debiasing strategy, it can be seamlessly incorporated into various SGG architectures [^3] Following the mainstream and concurrent unbiased SGG works, we also only focus on two-stage frameworks. As for one-stage models, a similar idea can be applied at the image-level, and we leave it for future works. and consistently improve their performance. Unsurprisingly, CFA can achieve a new state-of-the-art performance on the trade-off between different metrics. Extensive ablations and results on multiple SGG tasks and backbones have shown the generalization ability and effectiveness of CFA.
In summary, we make three contributions in this paper:
@@
    We reveal the issue of existing re-balancing methods, i.e., the lack of triplet feature diversity of tail categories. To this end, we are the first to tackle unbiased SGG from the perspective of increasing the diversity of triplet features.
@@
    We propose the model-agnostic CFA for unbiased SGG, which is an efficient and novel compositional learning framework that spans the feature space of the tail categories by two independent plug-and-play modules.
@@
    Extensive results show the effectiveness of CFA, i.e., it achieves a new SOTA performance on SGG benchmarks.
@@ 2 Related Work
Unbiased Scene Graph Generation. Biased predictions prevent further use of scene graphs in real-world applications. Recent unbiased SGG works can be roughly divided into three main categories: 1) Re-balancing: It alleviates the negative impact of long-tailed predicate distribution by re-weighting or re-sampling [41, 27, 31, 26, 8]. 2) Unbiased Inference: It makes unbiased predictions based on biased models [35, 42]. 3) Noisy Label Learning: It reformulates SGG as a noisy label learning problem and corrects these noisy samples [21, 22]. In this work, we point out the drawbacks of existing re-balancing methods and study unbiased SGG from the new perspective of feature augmentation.
Feature Augmentation. Data augmentation is a prevalent training trick to improve models’ performance. Conventional data augmentation methods [9, 48] usually synthesize new samples in a hand-crafted manner. Compared to these image-level data augmentation methods, feature augmentation is another efficient way to improve models’ generalizability by directly synthesizing samples in the feature space [6, 25]. Compared to existing data augmentation applications, SGG is a sophisticated task that involves intrinsic and extrinsic features. In this work, we propose CFA to enrich the diversity of relation triplet features for debiasing.
Compositional Learning (CL). CL has been successfully applied to various computer vision tasks. As for visual scene understanding, some Human Object Interaction detection works [14, 18, 15] compose new interaction samples that significantly benefit both low-shot and zero-shot settings. To the best of our knowledge, only two unpublished SGG work [13, 19] also uses CL. Compared to [13], we have several key differences: 1) They aim to generate new representations which are close to the original triplet. Instead, we try to increase the diversity of triplet features. 2) They only change the entities and ignore the extrinsic features. 3) Their augmentation strategies are mainly based on the spatial locations or IoU of the entities. The second work [19] primarily addresses classification tasks. Our work differs from it in two aspects: 1) CFA increases the diversity of tail predicate features by leveraging both intrinsic and extrinsic information. 2) we aim to improve robustness and performance under long-tailed predicate distributions.
Figure 3: The illustration of unbiased SGG framework with CFA.
Figure 4: The pipeline of Intrinsic CFA (a) and Extrinsic CFA (b). The blue and green boxes represent operations on the query triplet features in Intrinsic-CFA and context triplet features in Extrinsic-CFA, respectively.
3 Approach
@@
In this section, we first revisit the two-stage SGG baselines in Sec. 3.1. Then, we detailedly introduce our CFA in Sec. 3.2, including intrinsic-/extrinsic-CFA (c.f. Figure 3). Finally, we introduce the training objectives in Sec. 3.3.
@@ 3.1 Revisiting the Two-Stage SGG Baselines
Since the mainstream SGG methods are two-stage models, we review the two-stage SGG framework3 here [44, 36]. A typical two-stage SGG model involves three steps: proposal generation, entity classification, and predicate classification. Thus, the SGG task $P(\mathcal{G}|I)$ is decomposed into:
@@
Proposal Generation $P(\mathcal{B}|I)$. This step aims to generate all bbox proposals $\mathcal{B}$. Given an image $I$, they first utilize an off-the-shelf object detector (e.g., Faster R-CNN [34]) to detect all the proposals $\mathcal{B}$ and their visual features $\{\bm{v}_{i}\}$.
Entity Classification $P(\mathcal{O}|\mathcal{B},I)$. This step mainly predicts the entity category of each $\bm{b}_{i}$ $\in$ $\mathcal{B}$. Given a visual feature $\bm{v}_{i}$ and proposal $\bm{b}_{i}$, they use an object context encoder $\mathtt{Enc}_{obj}$ to extract the contextual entity representation $\bm{f}_{i}$:
@@
where $\bm{w}_{i}$ is the GloVe embedding [33] of predicted $\widehat{o}_{i}$. It is worth noting that both $\mathtt{Enc}_{cls}$ and $\mathtt{Enc}_{rel}$ often adopt a sequence model (e.g., Bi-LSTM [44], Tree-LSTM [36], or Transformer [42]) to better capture context. After relation feature encoding, they use a relation classifier $\mathtt{Cls}_{rel}$ to predict the relation $\widehat{r}_{ij}$ between any two subject-object pairs:
@@
3.2 CFA: Compositional Feature Augmentation
In this paper, we treat a relation triplet feature2 in SGG as a combination of two components: intrinsic feature and extrinsic feature, which refer to the intrinsic information (i.e., the subject and object itself) and extrinsic information (i.e., the contextual objects and stuff), respectively. Correspondingly, our CFA consists of intrinsic-CFA and extrinsic-CFA to augment these two types of features. For ease of presentation, we call the “targeted tail predicate triplet whose feature to be enhanced” as “query triplet”, and the “relation triplet whose image is used to provide context” as “context triplet”. Besides, to facilitate the feature augmentation, we store all the visual features of all tail predicate triplets (i.e., visual features $\bm{v}$ of two entities and their union feature $\bm{u}$) in a feature bank before training. Besides, we adopt the repeat factor $\eta$ = $\max(1,\eta_{r})$ [12, 26] to sample images to provide enough context triplet and query triplet for augmentation, where $\eta_{r}$ = $\sqrt{\lambda/{f_{r}}}$, $f_{r}$ is the frequency of predicate category $r$ on the entire dataset, and $\lambda$ is hyperparameter. The unbiased SGG pipeline with CFA is shown in Figure 3.